OpenStudy (susanna):

Evaluate cos (Tan^-1 squareroot of 3/3 + sin^-1 1/2). I think it's 1/2.

OpenStudy (susanna):

cos(Arctan (square root of 3)/3 + Arcsin 1/2)

OpenStudy (susanna):

@mathstudent55

OpenStudy (susanna):

@Michele_Laino

OpenStudy (mathstudent55):

Is this the problem? \(\cos [\tan^{-1} (\dfrac{\sqrt 3}{3}) + \sin^{-1} \dfrac{1}{2}]\)

OpenStudy (susanna):

yes

OpenStudy (mathstudent55):

As usual, we do this one step at a time. We start with the arctan and the arcsin. \(=\cos [30^o + 30^o]\) \(= \cos 60^o\) \(= \dfrac{1}{2} \) You are correct.

OpenStudy (susanna):

thanx

OpenStudy (mathstudent55):

yw